9. An ionized oxygen molecule (O2) at point A has charge t and moves at 2.00 x 103 m/s in the positive x-direction. A constant electric force in the negative x-direction slows the
molecule to a stop at point B, a distance of 0.750 mm pastA on the x-axis. Calculate (a) the x-component of the electric field and (b) the potential difference between points A and B.
Answer #1

using work energy theorem

e E d = 0.5 m v^2

1.6* 10^-19* E* 0.75* 10^-3 = 0.5* 16* 1.67* 10^-27* 2000^2

E = 445.33 N/C

======

b)

V = Ed = 445.33* 0.75* 10^3

V = 0.334 V